Opened 5 years ago

Closed 3 years ago

#499 closed flight processing (fixed)

IG13/11, flight day 224a/2013, Isafjord

Reported by: knpa Owned by: benj
Priority: alpha 4 low Milestone: 2013 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tipo)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/greenland_iceland/IG13_11-2013_224a_Isafjord

Data arrived from ARSF via network transfer on 23/08/2013

IG13/11 Scientific objective: unknown

IG13/11 Priority: 7

PI: Susan Conway

Sensors:

Camera (Dispatched 18/10/2013)
Hawk (Dispatched 18/10/2013)
Eagle (Dispatched 18/10/2013)
LiDAR (Dispatched 18/10/2013)

Change History (18)

comment:1 Changed 5 years ago by tipo

Processed nav but couldn't get a very good solution - will need to be redone when we get better basestation data.
Processed Hyperspectral and found SCTs. Will need to be double checked when the nav is redone.

Line Eagle Hawk
1 -0.01 -0.01
2 -0.01 -0.01
3 -0.01 -0.11
4 -0.01 -0.01
5 -0.01 -0.01
6 -0.01 -0.01
7 -0.01 -0.01
8 -0.01 -0.01
9 -0.01 -0.01
10 -0.01 1.5 (no nav)
11 -0.01 -0.01
12 1.40 (no nav) -0.01
13 -0.01 -0.01
14 3.05 (no nav) -0.01
15 -0.01 -0.01
16 -0.01 -0.01
17 -0.01 -0.01
18 -0.01 -0.01
19 -0.01 -0.01
20 -0.01 -0.01
21 -0.01 -0.01
22 -0.01 -0.01
23 -0.01 -0.01
Last edited 5 years ago by tipo (previous) (diff)

comment:2 Changed 5 years ago by knpa

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 low

comment:3 Changed 5 years ago by tipo

Created RCD Hyperspectral and LiDAR deliveries.

comment:4 Changed 5 years ago by tipo

Hyperspectral processed with LiDAR patched DEM as the ASTER only one resulted in large distortions where the height changed quickly, ie. the sides of the fjords

comment:5 Changed 5 years ago by emca

Delivery checked all sensors. Was the nav reprocessed with a better base station? If so, please update the ticket to reflect this. If not then a note needs to be added to the readme to mention some uncertainty in the data. The handwritten log sheets contain some info that can be added to the relevant notes section of the pdf readme:

  • weather conditions
  • problems with RCD
  • LiDAR data with low returns

LiDAR delivery check:
Extra xml file in dem directory
Navigation file name should be renamed to be consistent with naming structure (PROJCODE-YEAR_JULDAY)

RCD Delivery Check:
No issues, ready for delivery

Hyperspectral Delivery Check:
DEM needs to have ASTER added to its name as its a combined LiDAR aster DEM.
Didn't get time to check all the level1 files, just Eagle 1-8 and Hawk 16-23. Or to check the tiff files in tmp_cmd_check. All other checks are complete. You can get someone else to finish checking the level1 and tif files this week files or I will do so on my return next week.

comment:6 Changed 5 years ago by emca

Finished Hyperspectral delivery check. Level 1 files all look okay. No issues with the tiff images.
Should mention in the readme that there are underflows (particularly in the Hawk) resulting in gaps in the data.

For the camera KML file, would be better if you could remove the (0,0) coordinates from the non-tagged images as they are showing up off the south coast of Africa in google earth.

comment:7 Changed 5 years ago by tipo

Remade logsheet, renamed navigation files (Are we sure this is correct? proj_tidy doesn't recognise them with the changed names.)

comment:8 Changed 5 years ago by tipo

Remade hyperspectral readme.

comment:9 Changed 5 years ago by tipo

RCD delivery remade using minimal tagging as the few images that could be tagged were all being placed at the same position.

comment:10 Changed 5 years ago by emca

RCD: Remove the Image Collection "overview of you photographs" bit as there are no images displayed. Also add something to say that the KML file only contains the Eagle/Hawk and LiDAR but no camera data due the issues during data collection.

Lidar: Nav files has - instead of _ which is getting picked up by proj_tidy as incorrect

Hyperspectral: DEM is being picked up as incorrect by proj_tidy as it has lidar in the name (proj_tidy should be updated to accept this). Need to change the dem description in *.hdr to match the dem file name, and also update the DEM name in the readme sample command.

The ipas_honeywell folder has a load of dodgy named files, some of which may not be needed. Only keep those which were used for the processing. Any that are needed should be renamed as appropriate.

comment:11 Changed 5 years ago by tipo

Fixed the DEM header.

Updated the example commands.

Remade the RCD and Hyperspectral readmes. (Also added an explanation of the LiDAR patched DEM in the DEM section of the Hyperspectral readme.)
Renamed nav files.

The extra files in the ipas_honeywell directory are from the network adjustment I did on the basestations. I've moved them to the basestation directory for now rather than deleting them entirely.

Last edited 5 years ago by tipo (previous) (diff)

comment:12 Changed 5 years ago by emca

Data is all good and ready for delivery.

proj_tidy is still complaining about the basestation files (because of the naughty characters in the filenames).

comment:13 Changed 5 years ago by emca

Zipped the hyperspectral mapped files

comment:14 Changed 5 years ago by tipo

Dispatched deliveries 18/10/2013

comment:15 Changed 5 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Description modified (diff)
  • Owner set to benj

comment:16 Changed 5 years ago by tipo

Aproximate lat & lon for the RCD images found by comparing timestamps with the nav file. The RCD's clock doesn't sync up exactly with the nav's one, (it's roughly half an hour slow), so matching them exactly isn't possible, but plotting the coords in Google Earth, i'm confident they're pretty close - within a couple of hundred metres.
Created a delivery with the images(including the ones previously discarded for being mistakenly identified as out of bounds) tagged with the new coords.

comment:17 Changed 4 years ago by tec

Archival
Starting

comment:18 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data available from NEODC.

Note: See TracTickets for help on using tickets.